On H-closed spaces

M. N. Mukherjee and S. Raychaudhuri, Calcutta, India

Keywords: H-closed space, almost regular space, θ-adherence, θ-complete adherence, θ-cover, lower and upper θ^{∗}-continuous function AMS Subject classification: 54D20, 54D25

Abstract: In this article a study of H-closed topological spaces from a new stand point is undertaken. The results, thus obtained are applied to derive certain interesting techniques which ultimately lead us to the characterizations of such a space in the presence of almost regularity. Finally a fixed set theorem in analogy to that for compact spaces, is obtained for an H-closed space.